Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (g.Maccaferri at Sat Feb 20 08:19:12 MET 1999) Success

HartRAO (Jonathan Quick at Sat Feb 20 08:24:49 MET 1999) Success

Shanghai (Huang Xinyong at Mon Feb 22 10:59:24 MET 1999) Success
It is fine.

Effelsberg (dgraham at Mon Feb 22 15:35:35 MET 1999) Minor failures
video converter 11 has problems around 2112

Westerbork (Tony Foley at Mon Feb 22 21:26:33 MET 1999) Success
No known problems, but note that the tape unit was changed and the head parameters recalibrated during day 50.

Noto (C.Trigilio at Mon Feb 22 23:34:03 MET 1999) Minor failures
Lost the first two scans because of a problem at the tape. Some minut lost (from 20:38 to 20:45, source DA193) for antenna in prelimit.

Urumqi (Liu Xiang at Tue Feb 23 05:10:52 MET 1999) Minor failures
It is clear. The clock offset is 1.14us at the end. From UT 01:31:00 TO 03:18:04 the data is invalid because of the MK3 problem.

Medicina (A. Polatidis at Tue Feb 23 18:08:19 MET 1999) Success
No known problems

Jodrell (Alastair Gunn at Wed Feb 24 12:47:39 MET 1999) Success
No known problems.

Cambridge (Alastair Gunn at Wed Feb 24 12:48:04 MET 1999) Success
No known problems.

Noto (C.Trigilio at Fri Feb 26 11:36:35 MET 1999) Minor failures
BBC12 not locked to the right frequency

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:17:50 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: eh005.antab. Read the comments in the file to get further information. (Missed Ur log file)